# UniAudio 2.0: A Unified Audio Language Model with Text-Aligned Factorized Audio Tokenization

### v2

@@ 5.4 The Performance on Seen Tasks
In this section, we evaluate UniAudio 2.0 on seen pre-training tasks spanning speech, general audio, and music/song. We follow standard benchmarks and protocols and compare against both unified audio/speech foundation models (e.g., MiMo-Audio and Qwen2.5-Omni) and strong task-specific systems (e.g., MusicGen), depending on task availability. As summarized in Table 4, we can see that (1) UniAudio 2.0 obtains good performance in both speech generation and recognition tasks (e.g. ASR, TTS, Instructed TTS). Moreover, UniAudio 2.0 supports multiple languages, including English, Mandarin Chinese, and Cantonese (see Appendix B.6); (2) For general audio, UniAudio 2.0 also achieves performance that is competitive with previous SOTA task-specific models (e.g. Stable Audio); (3) For the music modality, it can effectively complete music/song understanding and generation tasks. Overall, UniAudio 2.0 consistently performs strongly across multiple benchmarks.
5.5 The Performance on Few-shot Tasks
In this part, we investigate the generalization ability of UniAudio 2.0 to unseen tasks. Following the setting of GPT-3 (Mann et al., 2020), we conduct a series of few-shot tasks covering both generation and understanding. For the few-shot generation tasks, we follow (Yang et al., 2024a; Zhang et al., 2025) and conduct few-shot speech denoising and voice conversion. For the few-shot understanding tasks, we conduct few-shot sound classification and emotion classification. Table 5 shows the experimental results: UniAudio 2.0 achieves better generalization on most benchmarks, especially in the 1-shot setting.
@@ 5.5 The Performance on Few-shot Tasks
5.6 The Performance on Zero-shot Tasks
In this part, we further conduct more challenging zero-shot tasks to evaluate the model’s generalization ability to unseen tasks. Specifically, we first evaluate the text capability on the MMLU benchmark (Hendrycks et al., 2020) in a zero-shot setting. Following (Fang et al., 2024), we use the InstructS2S-Eval benchmark to evaluate speech conversation ability in a zero-shot setting. Lastly, we design three tasks: dysarthric speech recognition (DSR) (Kim et al., 2008), speech-sound generation (Speech-S), and audio-prompt- and caption-guided speech generation (A-I-TTS). Table 6 shows the results; we have the following observations: (1) UniAudio 2.0 has strong text understanding ability, and introducing audio modalities does not significantly degrade text performance; (2) it performs well on many unseen tasks, e.g., speech-to-text conversation and dysarthric speech recognition; (3) it shows some generalization ability to newly defined tasks. For example, for speech-sound generation, we ask the model to generate speech and sound events based on the text content and sound tags. Furthermore, we observe that the model can follow instructions to use an audio prompt’s timbre and a caption’s style to synthesize the desired speech. However, these tasks are not solved perfectly, and we summarize the limitations in Appendix C.
@@ 5.7 Ablation Study
The influence of reasoning tokens We first explore the influence of reasoning tokens. As Table 2 and 3 show, when we do not use reasoning tokens in both understanding and generation settings, the performance is worse than combining reasoning tokens with reconstruction tokens. In Appendix A.3, we present more ablation studies about ReasoningCodec, including the influence of GRPO training loss A.3.1, the influence of multi-expert semantic encoders A.3.2, the effectiveness of FiLM A.3.3, the comparision with previous semantic tokenizer A.3.6, and the influence of classifier-free guidance (CFG) for the reconstruction performance A.3.7.
The influence of multi-stage training We compare performance across different training stages. As Table 7 shows, removing Stage 4 (the mid-training stage) degrades performance on several tasks, especially those related to text capability. We speculate that Stage 4 provides additional text exposure and longer-context training that help preserve and strengthen text understanding. Furthermore, without this mid-training stage, the model performs poorly on few-shot and zero-shot evaluations. These results suggest that increasing data and task diversity (by introducing auditory sentences) is key to improving generalization to unseen tasks.
The influence of layer specialization We conduct experiments to explore the effectiveness of the proposed functional layer specialization. Specifically, we drop the audio understanding and generation experts and directly train the cross-modal expert with Stage 3. As Table 7 shows, we find that the model’s performance significantly decreases on multiple benchmarks. This phenomenon further validates the effectiveness of our proposed unified AR structure.
The influence of model size We explore the influence of the cross-modal expert’s model size. Due to resource limitations, we use LLaMA 3.2 1B and LLaMA 3.2 3B as the cross-modal expert, respectively. For the 1B variant, we use the same training strategy as the 3B version. Table 7 shows the results: compared with the 3B model, the 1B model’s performance significantly decreases in multiple bencharks. In Appendix B.6.5, we also show that its generalization ability on few-shot and zero-shot tasks is much weaker. We believe model size is a key factor that determines the limit of model capability, especially for multi-task foundation models. Similarly, OpusLM (Tian et al., 2025a) also found that model size has a strong influence on unified understanding and generation models.
@@ 6 Conclusion
In this study, we investigate how to build a unified audio foundation model that supports both understanding and generation. We propose ReasoningCodec, which factorizes audio into reasoning tokens and reconstruction tokens, and train UniAudio 2.0 with a unified autoregressive architecture and a multi-stage, multi-task training strategy. Experiments show strong performance on seen speech, sound, and music tasks, as well as encouraging few-shot and zero-shot generalization to unseen tasks. Extensive ablation studies suggest that scaling data/task diversity and model size is key to improving generalization on unseen tasks. In the future, we plan to scale both the model and the training data to further improve generalization. We discuss limitations in Appendix C.
@@ A.2.2 Stage 2: GRPO for Reasoning Branch
A.2.3 Stage 3: Reconstruction waveform by flow-based diffusion decoder
In this stage, we freeze the reasoning branch and train the reconstruction branch to reconstruct the waveform. Specifically, we first use a multi-expert semantic encoder to extract semantically rich features for speech, music, and sound. Then, we use three parallel VQ groups to quantize these features into discrete tokens, with a 1:1:6 allocation of VQ layers across the three groups. We use six VQ layers for the last group because we expect it to encode more acoustic details for better reconstruction. Together, the three groups produce 8 discrete tokens per frame, which we refer to as reconstruction tokens. Finally, we follow SimpleSpeech (Yang et al., 2024b) and build a flow-based scalar diffusion decoder to recover the waveform from these reconstruction tokens.
We do not directly condition waveform reconstruction on reasoning tokens for two reasons: (1) the reconstruction tokens already contain the necessary information provided by reasoning tokens, and (2) conditioning would increase the sequence length for the diffusion transformer. In early experiments, upsampling reasoning tokens and combining them with reconstruction tokens brought only marginal improvements. Instead, we use FiLM (Perez et al., 2018) to inject reasoning-token information into the multi-expert semantic encoder. Empirically, we observe no significant degradation in reconstruction quality.
For SQ-Codec, we follow the training recipe of SimpleSpeech (Yang et al., 2024c) and set the latent dimension to 136. We train SQ-Codec on the same dataset used in Stage 3 of Table 10.
@@ A.3.2 The effectiveness of multi-expert semantic features
Different audio modalities may emphasize different types of semantic information. We therefore use multiple expert encoders to extract modality-specific semantic features and apply separate VQ heads to quantize them. To demonstrate the effectiveness of this strategy, we build two baselines: (1) concatenating all semantic features and applying query-based quantization; and (2) concatenating all semantic features and applying standard RVQ. Table 16 shows the results. From the reconstruction perspective, query-based quantization performs better, while group-wise VQ yields better token modeling performance, especially for music. We attribute this to using separate VQ layers for different semantic features. Although query-based quantization can further improve reconstruction, it requires multiple transformer encoders. Considering the additional inference cost, we do not apply query-based quantization in the reconstruction branch.
@@ A.3.4 The subjective evaluation for audio tokenizer
MimiCodec (8 RVQ) (Défossez et al., 2024) 12.5/100 2048 86.7 $\pm$ 2.1 - -
ReasoningCodec 12.5/100 1024 90.5 $\pm$ 2.8 80.8 $\pm$ 2.0 86.6 $\pm$ 2.3
Table 14 shows the subjective evaluation. We can see that our proposed ReasoningCodec obtains best performance on both speech, sound, and music reconstruction.
@@ A.3.7 The influence of classifier-free guidance and decoding steps
Table 16: Comparison of different quantization strategies.
@@ B.3 The details of multiple stage training
Figure 5: Pre-training stages. For simplicity, the audio generation expert layers includes both generation expert and local decoder layers.
@@ B.6.1 ASR performance comparison
Mimo-Audio-Instruct-7B 3.50 35.43 29.81 7.01
Qwen2.5-Omni-7B 3.92 5.52 1.3 2.89
Step-Audio-chat-3B 3.11 8.44
Whisper-large-v3 1.81 3.55 6.8 1.47
@@
Table 20: ASR results on WenetSpeech-Yue Benchmark.
@@
Whisper-large-v3 36.8 31.5
SenseVoice-Yue 15.7 6.0
Qwen2.5-Omni-7B 23.5 31.0
@@ B.6.2 TTS performance comparison
Table 21: TTS results on Seed-TTS (EN/ZH), LibriSpeech-clean, and Cantonese TTS. We report the WER and DNS-MOS score (WER / DNS-MOS).
@@
Table 22: Ablation studies for multi-stage training, We include additional tasks that are not reported in Table 7. For Few-shot VC and Few-shot Sound, we report results under the 1-shot setting. We use NA to indicate that the model cannot perform the task without the stage 4 training.
@@ B.6.3 Audio/music caption performance comparison
SALMON 13B 0.355 5.41 8.06 3.24 7.92
Music Flamingo 4.74 8.19
Audio Flamingo 0.51 5.25 8.01
Audio Flamingo 2 0.58 5.71 8.13
Audio Flamingo 3 0.79 6.34 8.24 5.77 8.24
Qwen2.5-Omni 0.39 5.61 8.21 5.33 7.90
@@ B.6.4 Audio generation performance comparison
AudioLDM2-large 1.57 170.31 0.41
Stable Audio Open 2.14 78.24 0.29
AudioGen 1.42 186.53 0.45
Ours 3.26 50.69 0.17
@@ B.6.5 Zero-shot experiments for 1B model
Text LLAMA 3.2 1B 34.14
LLAMA 3.2 3B 47.63
@@
S2S LLAMA-Omni 3.47 / 3.99
SpeechGPT 2.19 / 2.98

### v3

@@ 5.4 The Performance on Seen Tasks
In this section, we evaluate UniAudio 2.0 on tasks seen during pre-training, spanning speech, general audio, and music/song domains. We follow standard benchmarks and evaluation protocols, and compare against both unified audio–speech foundation models (e.g., MiMo-Audio (Zhang et al., 2025) and Qwen2.5-Omni-7B (Xu et al., 2025)) and strong task-specific systems (e.g., MusicGen (Copet et al., 2023)), where applicable. [^3] Although many tasks have strong task-specific baselines, we adopt unified base models to enable fair and consistent comparisons across tasks.
As summarized in Table 5, we observe that: (1) UniAudio 2.0 achieves strong performance on both speech generation and recognition tasks (e.g., ASR, TTS, and instructed TTS), and supports multiple languages, including English, Mandarin Chinese, and Cantonese (see Appendix B.6). Compared with the previous state-of-the-art speech foundation model, MiMo-Audio 7B, UniAudio 2.0 achieves better performance on both ASR and TTS tasks with only 3B parameters; (2) for general audio, UniAudio 2.0 remains competitive with previous state-of-the-art task-specific models (e.g., Stable Audio), and also outperforms Qwen2.5-Omni 7B on audio captioning tasks; and (3) for the music modality, it effectively supports both music/song understanding and generation tasks. Overall, UniAudio 2.0 consistently performs well across multiple benchmarks.
5.5 Performance on Few-shot Tasks
In this section, we investigate the generalization ability of UniAudio 2.0 to unseen tasks. Following the setting of GPT-3 (Mann et al., 2020), we evaluate our model on a series of few-shot tasks covering both generation and understanding.
For few-shot generation, we follow (Yang et al., 2024a; Zhang et al., 2025) and evaluate speech denoising and voice conversion. For few-shot understanding, we consider sound classification and emotion classification. Table 6 summarizes the experimental results, showing that UniAudio 2.0 achieves strong generalization performance on most benchmarks, especially in the 1-shot setting.
@@ 5.5 Performance on Few-shot Tasks
5.6 Performance on Zero-shot Tasks
We further evaluate UniAudio 2.0 on more challenging zero-shot tasks to assess its generalization ability to unseen scenarios. Specifically, we first evaluate its text understanding capability on the MMLU benchmark (Hendrycks et al., 2020) in a zero-shot setting. Following (Fang et al., 2024), we use the InstructS2S-Eval benchmark to evaluate speech conversation ability without task-specific examples. In addition, we design three new tasks: dysarthric speech recognition (DSR) (Kim et al., 2008), speech-sound generation (Speech-S), and audio-prompt- and caption-guided speech generation (A-I-TTS). Table 7 reports the results. We make the following observations. First, UniAudio 2.0 demonstrates strong text understanding ability, and the introduction of audio modalities does not significantly degrade text performance. Second, the model performs well on many unseen tasks, such as speech-to-text conversation and dysarthric speech recognition. Third, it exhibits promising generalization ability on newly designed tasks. For example, in speech-sound generation, the model is required to generate speech and sound events based on textual content and sound tags. Moreover, we observe that UniAudio 2.0 can follow instructions to leverage the timbre of an audio prompt and the style specified in a caption to synthesize the desired speech. Nevertheless, these tasks are not yet solved perfectly, and we summarize the limitations in Appendix C.
@@ 5.7 Ablation Study
Influence of reasoning tokens. We first investigate the impact of reasoning tokens. As shown in Tables 2 and 3, removing reasoning tokens in both understanding and generation settings leads to worse performance than combining reasoning and reconstruction tokens. In Appendix A.3, we present additional ablation studies on ReasoningCodec, including the influence of the GRPO training loss (Appendix A.3.1), the effect of multi-expert semantic encoders (Appendix A.3.2), the effectiveness of FiLM (Appendix A.3.3), comparisons with previous semantic tokenizers (Appendix A.3.6), and the impact of classifier-free guidance (CFG) on reconstruction performance (Appendix A.3.7).
Influence of multi-stage training. We compare performance across different training stages. As shown in Table 8, removing Stage 4 (the mid-training stage) degrades performance on several tasks, particularly those related to text capability. We attribute this degradation to the reduced text exposure and shorter context length during training, which weakens text understanding. Furthermore, without this mid-training stage, the model performs poorly on few-shot and zero-shot evaluations. These results suggest that increasing data and task diversity, for example by introducing auditory sentences, is crucial for improving generalization to unseen tasks.
Influence of layer specialization. We conduct experiments to evaluate the effectiveness of the proposed functional layer specialization. Specifically, we ablate the audio understanding and generation experts and train only the cross-modal expert using Stage 3. As shown in Table 8, this modification leads to a significant performance degradation across multiple benchmarks. These results further validate the effectiveness of our unified autoregressive architecture. Moreover, our architectural modifications introduce no additional infrastructure complexity and remain fully compatible with tensor, pipeline, and context parallelism.
Influence of model size. We investigate the impact of the cross-modal expert’s model size. Due to resource constraints, we use LLaMA 3.2 1B and LLaMA 3.2 3B as the cross-modal expert, respectively, and adopt the same training strategy for both variants. As shown in Table 8, the 1B model exhibits a significant performance drop compared to the 3B model across multiple benchmarks. Appendix B.6.5 further shows that its generalization ability on few-shot and zero-shot tasks is substantially weaker. These results indicate that model size is a critical factor in determining model capacity, especially for multi-task foundation models. Consistent with our findings, OpusLM (Tian et al., 2025a) also reports that model scale plays a crucial role in unified understanding and generation models.
@@ 6 Conclusion
In this work, we present a unified audio foundation model that supports both understanding and generation. We propose ReasoningCodec, which factorizes audio into reasoning and reconstruction tokens, and train UniAudio 2.0 using a unified autoregressive architecture together with a multi-stage, multi-task training strategy. Extensive experiments demonstrate strong and consistent performance on seen speech, sound, and music tasks, as well as promising few-shot and zero-shot generalization to unseen scenarios.
Our ablation studies further indicate that scaling data and task diversity, as well as model capacity, plays a crucial role in improving generalization. In future work, we will continue to scale both model size and training data to further enhance the robustness and generalization ability of unified audio foundation models. The limitations of our approach are discussed in Appendix C.
@@ A.2.2 Stage 2: GRPO for Reasoning Branch
A.2.3 Stage 3: Waveform Reconstruction with a Flow-based Diffusion Decoder
In this stage, we freeze the reasoning branch and train the reconstruction branch to recover waveforms from discrete tokens. Specifically, we first employ a multi-expert semantic encoder to extract semantically rich features from speech, music, and sound signals. These features are then quantized into discrete tokens using three parallel VQ groups, with a 1:1:6 allocation of VQ layers across the three groups. We assign six VQ layers to the last group to encode finer-grained acoustic details for high-fidelity reconstruction.
Overall, the three groups produce eight discrete tokens per frame, which we refer to as reconstruction tokens. Following SimpleSpeech (Yang et al., 2024b), we further construct a flow-based scalar diffusion decoder to reconstruct waveforms from these tokens.
We do not directly condition waveform reconstruction on reasoning tokens for two main reasons. First, reconstruction tokens already capture the essential information provided by reasoning tokens. Second, explicit conditioning would significantly increase the sequence length of the diffusion transformer. In preliminary experiments, upsampling reasoning tokens and concatenating them with reconstruction tokens resulted in only marginal improvements. Instead, we adopt FiLM (Perez et al., 2018) to inject reasoning-token information into the multi-expert semantic encoder. Empirically, this design introduces no noticeable degradation in reconstruction quality.
For SQ-Codec, we follow the training protocol of SimpleSpeech (Yang et al., 2024c) and set the latent dimension to 136. The model is trained on the same dataset used in Stage 3, as summarized in Table 11.
@@ A.3.2 The effectiveness of multi-expert semantic features
Different audio modalities may emphasize different types of semantic information. We therefore use multiple expert encoders to extract modality-specific semantic features and apply separate VQ heads to quantize them. To demonstrate the effectiveness of this strategy, we build two baselines: (1) concatenating all semantic features and applying query-based quantization; and (2) concatenating all semantic features and applying standard RVQ. Table 16 shows the results. From the reconstruction perspective, query-based quantization performs better, while group-wise VQ yields better token modeling performance (semantic information), especially for music. We attribute this to using separate VQ layers for different semantic features. Although query-based quantization can further improve reconstruction, it requires multiple transformer encoders. Considering the additional inference cost, we do not apply query-based quantization in the reconstruction branch.
@@ A.3.4 The subjective evaluation for audio tokenizer
For the subjective evaluation, we follow previous works (Yang et al., 2025a; Kreuk et al., 2022) conduct the MUSHRA test. Table 4 shows the subjective evaluation. We can see that our proposed ReasoningCodec obtains best performance on both speech, sound, and music reconstruction.
@@ A.3.7 The influence of classifier-free guidance and decoding steps
Table 16: Comparison of different quantization strategies. For each strategy, we control the same VQ layers (8 VQ layers), training data, and decoder.
@@ B.3 The details of multiple stage training
Figure 5: Overview of the multi-stage pre-training strategy of UniAudio 2.0. The model is trained in a progressive manner, including an understanding warm-up stage, a generation warm-up stage, and a modality alignment stage, with different expert layers gradually unfrozen. Stream-wise fusion is employed to merge text and audio streams under a unified autoregressive objective. For simplicity, the audio generation expert layers include both the generation expert and the local decoder.
@@ B.6.1 ASR performance comparison
Mimo-Audio-Instruct-7B (Zhang et al., 2025) 3.50 35.43 29.81 7.01
Qwen2.5-Omni-7B (Xu et al., 2025) 3.92 5.52 1.3 2.89
Step-Audio-chat-3B (Huang et al., 2025) 3.11 8.44
Whisper-large-v3 (Radford et al., 2023) 1.81 3.55 6.8 1.47
@@
Table 20: ASR results on WenetSpeech-Yue ASR Benchmark.
@@
Whisper-large-v3 (Radford et al., 2023) 36.8 31.5
SenseVoice-Yue (Li et al., 2025b) 15.7 6.0
Qwen2.5-Omni-7B (Xu et al., 2025) 23.5 31.0
@@ B.6.2 TTS performance comparison
Table 21: TTS results on Seed-TTS (EN/ZH), LibriSpeech-clean, and Cantonese TTS. We report the WER and DNS-MOS score (WER / DNS-MOS). The CosyVoice-Yue is the official checkpoint from WenetSpeech-Yue (Li et al., 2025b).
@@
Table 22: Ablation studies for multi-stage training, We include additional tasks that are not reported in Table 8. For Few-shot VC and Few-shot Sound, we report results under the 1-shot setting. We use NA to indicate that the model cannot perform these few-shot task without the stage 4 training.
@@ B.6.3 Audio/music caption performance comparison
SALMON 13B (Tang et al., 2024) 0.355 5.41 8.06 3.24 7.92
Music Flamingo (Ghosh et al., 2025a) 4.74 8.19
Audio Flamingo (Kong et al., 2024) 0.51 5.25 8.01
Audio Flamingo 2 (Ghosh et al., 2025b) 0.58 5.71 8.13
Audio Flamingo 3 (Goel et al., 2025) 0.79 6.34 8.24 5.77 8.24
Qwen2.5-Omni-7B (Xu et al., 2025) 0.39 5.61 8.21 5.33 7.90
@@ B.6.4 Audio generation performance comparison
AudioLDM2-large (Liu et al., 2023) 1.57 170.31 0.41
Stable Audio Open (Evans et al., 2025) 2.14 78.24 0.29
AudioGen (Kreuk et al., 2022) 1.42 186.53 0.45
UniAudio 2.0 (Ours) 3.26 50.69 0.17
@@ B.6.5 Zero-shot experiments for 1B model
Text LLAMA 3.2 1B (Dubey et al., 2024) 34.14
LLAMA 3.2 3B (Dubey et al., 2024) 47.63
@@
S2S LLAMA-Omni (Fang et al., 2024) 3.47 / 3.99
SpeechGPT (Zhang et al., 2023a) 2.19 / 2.98